What volume of each of the following solutions contains the indicated amount of dissolved solute?

40.0 g solute in 0.375 M NaF

Express your answer with the appropriate units.

40.0 g solute in 0.375 M CdCl2

Express your answer with the appropriate units.

10.0 g solute in 0.497 M K2CO3

Express your answer with the appropriate units.

10.0 g solute in 0.497 M Fe(ClO3)3

Express your answer with the appropriate units.



Concentrated acetic acid is 6 M.

What is the molarity of an HC2H3O2 solution prepared by diluting 70 mL of concentrated acid to a total volume of 400 mL?

Express your answer to two significant figures and include the appropriate units.



Concentrated nitric acid is 23 M.

What volume of HNO3 must be diluted with distilled water to prepare 6.00 L of 0.40 HNO3?

Express your answer to two significant figures and include the appropriate units.
1
Expert's answer
2020-12-10T07:50:20-0500

1. 40/42 x 0.375=2,51952633 L of NaF.

2. 40/183 x 0.375 = 0,58287796 L of CdCl2.

3. 10/138 x 0.497 = 0,14580235 L of K2CO3.

4. 10/306.5 x 0.497=0,0656467352 L of Fe(ClO3)3.

5. 0.07 x 6 /0.4 = 1,05 M. CH3COOH.

6. 6 x 0.4 /23 = 0,104347826 L of HNO3.